OpenStudy (anonymous):

Find the slope of the line that contains the points named. (Use fraction format.) R(0, 4), S(5, 0) m =

OpenStudy (anonymous):

= (0 - 4) / ( 5 - 0) = -4/5 (difference in y coordinates / difference in corresponding x coordinates)

OpenStudy (anonymous):

what about this 1

OpenStudy (anonymous):

X(2, -1), Y(0, 4) m =

OpenStudy (anonymous):

slope of line XY = (4-(-1)) / (0-2) = 5 /-2 = -5/2

OpenStudy (anonymous):

it doesnt matter which point you take first - it gives same result slope = ((-1)-4) / (2-0) = -5 / 2

OpenStudy (anonymous):

wow your smart my teachers says im in sChjool and we are checkinq over my friends answers

OpenStudy (anonymous):

what about C(3, 8), D(-2, 5) m =

OpenStudy (anonymous):

you should be able to do these now -ill start this one, for you to complete slope Of CD = (5 - 8) / (-2-3) = -3 /

OpenStudy (anonymous):

5

OpenStudy (anonymous):

no its -5: -2-3 = -5 so slope is -3/-5 = 3/5

OpenStudy (anonymous):

o ok thanks